OpenStudy (stylesprettygirl):

Can anyone help me with quarterly compounded interest any help would mean a lot need help asap questions in commments

OpenStudy (stylesprettygirl):

Thomas opened a savings account with an annual interest rate of 7% and an initial deposit of $5000. If his interest is compounded quarterly, how much is in Thomas’s account after 4 years? Round your answer to the nearest cent. interest compounded quarterly: A = P (1 +fraction numerator r over denominator 4 end fraction)4t a=5000(1+7/4) ^4x4

jimthompson5910 (jim_thompson5910):

The compound interest equation is \[\Large A = P\left(1+\frac{r}{n}\right)^{n*t}\]

jimthompson5910 (jim_thompson5910):

A = final amount after t years P = initial deposit r = interest rate in decimal form n = number of times the money is compounded per year t = number of years

jimthompson5910 (jim_thompson5910):

In this case A = unknown P = 5000 dollars r = 0.07 (7% = 7/100 = 0.07) n = 4 (quarterly ---> 4 times a year) t = 4 years

jimthompson5910 (jim_thompson5910):

\[\Large A = P\left(1+\frac{r}{n}\right)^{n*t}\] \[\Large A = 5000\left(1+\frac{0.07}{4}\right)^{4*4}\] \[\Large A = ??\] I'll let you compute the result

OpenStudy (stylesprettygirl):

828245 ^ is what I got thank you for the help if your still on am I right?

jimthompson5910 (jim_thompson5910):

that's way too big

jimthompson5910 (jim_thompson5910):

type `5000*(1+0.07/4)^(4*4)` into a calculator http://web2.0calc.com/

jimthompson5910 (jim_thompson5910):

let me know what you get @stylesprettygirl

OpenStudy (stylesprettygirl):

6599.65

jimthompson5910 (jim_thompson5910):

I got that as well

OpenStudy (stylesprettygirl):

Thank you soooooo much. Your the beeesssstttt

jimthompson5910 (jim_thompson5910):

you're welcome